OpenStudy (anonymous):

Given f(x) = 2x2 + 2x and g(x) = x + 1, find the function (f/g)(x).

OpenStudy (anonymous):

stick \(f\) on top of \(g\) to get \[\frac{f}{g}(x)=\frac{2x^2+2}{x+1}\] then maybe factor and cancel

OpenStudy (anonymous):

sorry typo, it is \[\frac{f}{g}(x)=\frac{2x^2+2x}{x+1}\]

OpenStudy (anonymous):

factor out the \(2x\) from the numerator to get \[\frac{f}{g}(x)=\frac{2x^2+2}{x+1}=\frac{2x(x+1)}{x+2}=2x\] if \(x\neq 1\)

OpenStudy (anonymous):

thank you, and good explaining :)